solve the differential equations
FE F EEL FREE
Answer #1

substitute y = Va » - V,1 + ༤ ཉི༔ = Vx2 + x de dx , butting these values in 0; v + x du X+ Vx ༡༩ - /z - x ( 1 + ༧) c( 1 ཀ༩) =x du tv. av 1-V I +V - V(-) 1-V = [+V-0 + 2004 1-ve = 1+ 2a lov » X do t+nga l-ve on deve & de Corted in the dev = = de Integ- Sings do - homo dv + Si de ► tan(v) - ava du = loll +C >> tan(a) – 5 An 117001 - holzel+G a tan( os) - Anf1+ v) 4a\ - kul